# A Set-Sequence Model for Time Series

### v1

@@ Appendix A Additional Literature Review
This approach naturally handles a variable number of units and allows immediate inference for unseen units because the same parameters are shared, but it ignores cross-unit dependencies. Representative methods include classical Vector Autoregressive (VAR) models [26], which scale quadratically in the cross-section, and a broad class of modern deep-learning time-series models: Transformer variants [41, 18, 42, 35, 20], state-space models [10, 39, 34], convolutional models [8], multilayer perceptrons [6, 37], and graph neural networks [36, 25].
@@
    Dynamic factor models. Linear versions [4, 29, 27, 28, 3] reduce dimensionality but require re-estimating loadings (and sometimes factors) whenever units change. Deep factor models [33] replace linear dynamics with RNNs, yet still demand refitting loadings and use local linear/Gaussian components that may miss nonlinear effects.
@@
Joint modelling is vital in finance—for instance, contagion in loan defaults [1, 2, 30]. Previous work often operated on one unit at a time [22, 14]. In portfolio optimisation, models typically impose a fixed ordering and limit cross-sectional width [15], overlooking exchangeability.
Appendix B Set-Seq model with Gated Selection
@@
where $\rho(\cdot)$ is an output projection. The gating, with $G_{i,j}\in[0,1]$ allows the summary statistic of unit i to mainly depend on the other units with ’similar’ features to unit i. We recover the standard Set-Seq model when we set $G_{i,j}=\frac{1}{M}$, for all $i,j$ where $M$ is the number of units. We provide experiments using the Gated Selection mechanism in Section 14.
@@
  • Most importantly, we only compute one G for all times for the sample, this has a large impact of the memory usage, removing the linear scaling with sequence length needed for gradient computations in the selection mechanism.
Appendix C Experiment Details
@@
Table 7: Hyper-parameter settings for each task.
Hyper-parameter Synthetic Mortgage Equity
Set-Seq layers 5 6 6
@@
All the experiments in the paper was conducted on a Linux cluster with 5 NVIDIA RTX A6000 GPUs, each with 49140 MB memory, running on CUDA Version 12.5. The cluster has 256 AMD EPYC 7763 64-Core Processor CPUs. For the synthetic task, each model was trained for less than 1 hour on 1 GPU. For the equities task, each train run (one run per validation year and random seed) took around 30 minutes on a single GPU. The experiments for the mortgage risk prediction task ran for around 1 hour per model training on one GPU.
@@
Appendix D Synthetic Task: Additional Details and Results
D.1 Ablations on the training input units sampling
@@ D.2 Interpretability
D.3 Kalman Filter Baseline
The synthetic task is described in Section 5. We assume that the transition matrix is known, and we also assume that the dynamics of $\lambda_{t}$, $\lambda_{t+1}=\beta\lambda_{t}+\alpha N_{t}$ are known, with the only unknown factor being not observing the true fraction of rare events $N_{t}$, but we only observe $\bar{N}_{t}$, where $\bar{N}_{t}$ is the observed fraction of rare events, only having access to n out of M units. The problem is now to infer the best estimate of $\lambda_{t}$ based on a partial observation of the fraction of rare events. We utilize a Kalman Filter [13] to estimate $\lambda_{t}$ with $\hat{\lambda}_{t}$. The resulting filter estimates
@@ D.3 Kalman Filter Baseline
[…]
@@
Appendix E Equity Portfolio Prediction: Additional Details and Results
Table 9: Out of sample annualized performance metrics net transaction costs 2002 - 2016, trained with a net sharpe ratio objective. Details of the models other than the Set-Sequence model are found in [11]. The training setting for the sequence models and Set-Sequence model is in Appendix C.
Model Net Sharpe Return % Std Dev %
@@ Appendix E Equity Portfolio Prediction: Additional Details and Results
We test the Set-Sequence model and the other baseline models under realistic transaction costs. In particular, for each day we subtract the transaction costs from the daily portfolio returns, which are used to compute the Sharpe ratio. Following [11], we use the following model for the transaction cost for portfolio weights $w_{t}$ at time t, and $w_{t-1}$ at t-1:
$\text{cost}(w_{t},w_{t-1})=0.0005\times||w_{t}-w_{t-1}||_{1}+0.0001\times||max(-w_{t},0)||_{1}.$
The return at time t, can then be expressed as $r_{net,t}=r_{t}-\text{cost}(w_{t},w_{t-1})$. We now train the Set-Seq model and the Sequence model baselines with a modified objective using the net return Sharpe as the loss function. In Table 9 we show the net Sharpe ratio after costs in Jan. 2002 - Dec. 2016. We see the gain from modeling the cross-section as the set-sequence model outperforms all the sequence models in terms of net Sharpe Ratio, as well as the domain specific CNN-Transformer.
@@
Table 10 shows the rank normalized features. In addition to a subset of the features in [5] we include features for daily return, weekly return, weekly volatility, daily volume. In order to keep the level information, we also include cross-sectional median features in Table 11, following [5].
Table 10: Firm-specific characteristics (six categories) used as features in the equity-portfolio-optimization task. Construction details are in the Internet Appendix of [5].
@@
r2_1 Short-term momentum A2ME Assets to market cap
r12_2 Momentum BEME Book-to-market ratio
r12_7 Intermediate momentum C Cash + short-term inv. / assets
r36_13 Long-term momentum CF Free cash-flow / book value
ST_Rev Short-term reversal CF2P Cash-flow / price
@@
Rel2High Closeness to 52-week high
@@
PM Profit margin RF Risk-free rate
@@
PCM Price-to-cost margin
Table 11: Cross-sectional median firm-characteristic variables and their stationary transformations ($t\text{Code}$).  Each feature is the cross-sectional median of the underlying firm characteristic. Note: The transformations ($t\text{Code}$) are (1) no transformation; (2) $\Delta x_{t}$; (3) $\Delta\log(x_{t})$; (4) $\Delta^{2}\log(x_{t})$.
Variable $t$Code Variable $t$Code Variable $t$Code Variable $t$Code
@@
Appendix F Mortgage Risk Prediction: Additional Details and Results
F.1 Additional Data Descriptions
We provide additional details about the CoreLogic Loan-Level Market Analytics Dataset. The dataset is over 870 GB and contains month-by-month transition data for mortgages in the US. The dataset starts in 1988, and has data to 2024, with data from 30000 ZIP codes in the US. We restrict ourselves to the top 4 ZIP codes, in terms of number of active loans, in the greater Los Angeles area and leave to follow up work to train on the full dataset. The ZIP Codes are: 92677, with 109642 loans in Orange County; 93065, with 98673 loans in Simi Valley; 91709, with 95497 loans in Chino Hills; 92336 with 94794 loans in Fontana. We follow the same data filtering procedure as in [22], and use a subset with 50 of the features included there, we filter out loans if of the following features are not present: FICO score, Original balance, Initial interest rate, and Current State. Some other features we use include Original LTV, Unemployment Rate, Current Interest Rate, and National Mortgage Rate. We deal with missing data in other features with the missing indicator method, see, for example [17]. After filtering $117,523$ loans remain, each active for on average 45 months, for a total of around 5 million loan month transitions. Figure 11 shows the empirical transition probabilities across the dataset, and Figure 11 shows the transition counts on the sampled test set. Figure 12 shows the number of active loans, the prepayment rate, and the foreclosure rate over time on the dataset. For example, note the elevated foreclosure rates during and after the 2008 financial crisis. Table 12 shows the full set of features for the Set-Sequence model.
@@ F.1 Additional Data Descriptions
Table 12: Feature definitions and possible values. We require that current state, FICO score, Original balance, and Initial Interest Rate are available, and use a missing indicator when other features are unavailable. The feature vector including the missing indicators has dimension 50.
@@
Paid Off, Foreclosure, Real-Estate Owned
@@
Time since origination $<1$ Year,  $<5$ Years,  $<10$ Years
Original term $<17$ Years
Number of 30-day delinquencies (past 12 months) 0–12
Number of 60-day delinquencies (past 12 months) 0–12
Number of 90+-day delinquencies (past 12 months) 0–12
Number of 30-day foreclosures (past 12 months) 0–12
@@ F.2 Model Fitting
F.3 Additional Experiments Mortgage Risk Prediction
F.3.1 Transition Analysis
Figure 13 shows the one vs rest AUC (see Section 5.2 for the definition) conditioned on the initial state for the Set-Sequence model, averaged over 10 seeds of samples on the test set, where for each seed we sample 25 sequences of 2500 loans from the test set.
@@ F.3.1 Transition Analysis
F.3.2 Year-by-year results
@@ Results
We compare the Set-Sequence model, the Gated Selection model, and the models from  [22] in the setting where we do a base training for 15 epochs up until year 2002, and then for each epoch the train-set gets extended one year further. We save a checkpoint each year the model is extended.
In addition we up-weight the more recent train data, this is described in Section F.3.2. Figure 15 and Figure 16 shows the results, comparing the Set-Seq, Gated Selection (see Appendix B), NN, and Logistic model, all using the same retraining method, and where we evaluate on one year at a time.
Figure 15: Average AUC, Cross-Entropy, and 30dd-60dd AUC for the Set-Seq, NN, and Logistic model, all using the same retraining method, and where we evaluate on one year at a time.
@@
We see that the challenging years around the financial crisis affects the logistic model the most. However, the relative performance of the Set-Seq model and the NN model remain broadly the same, although it widens somewhat during 2009.
F.3.3 Interpretability
@@
Appendix G Additional Metric Details
@@
$\operatorname{AUC}(k)=P(y_{1,k}>y_{2,k}|y_{1}=k,y_{2}\neq k)+\frac{1}{2}P(y_{1,k}=y_{2,k}|y_{1}=k,y_{2}\neq k),$

### v2

@@ Appendix B Additional Literature Review
This approach naturally handles a variable number of units and allows immediate inference for unseen units because the same parameters are shared, but it ignores cross-unit dependencies. Representative methods include classical Vector Autoregressive (VAR) models Sims (1980), which scale quadratically in the cross-section, and a broad class of modern deep-learning time-series models: Transformer variants Zhou et al. (2021); Liu et al. (2022); Zhou et al. (2022); Wu et al. (2021); Piao et al. (2024), state-space models Gu et al. (2022); Zhang et al. (2023); Wang et al. (2025b), convolutional models Fu et al. (2023), multilayer perceptrons Chen et al. (2023); Yi et al. (2023), and graph neural networks Wu et al. (2020a); Shang et al. (2021). Recent time series forecasting models Wang et al. (2025a; 2024b) are complementary in that they can be used as the sequence model backbone for the Set-Sequence model.
@@
    Dynamic factor models. Linear versions Chamberlain & Rothschild (1983); Stock & Watson (2002a; 2011; b); Bok et al. (2018) reduce dimensionality but require re-estimating loadings (and sometimes factors) whenever units change. Deep factor models Wang et al. (2019) replace linear dynamics with RNNs, yet still demand refitting loadings and use local linear/Gaussian components that may miss nonlinear effects.
@@
Joint modelling is vital in finance—for instance, contagion in loan defaults Anenberg & Kung (2014); Azizpour et al. (2018); Towe & Lawley (2013). Previous work often operated on one unit at a time Sadhwani et al. (2020); Khandani et al. (2010). In portfolio optimisation, models typically impose a fixed ordering and limit cross-sectional width Kisiel & Gorse (2023), overlooking exchangeability.
Appendix C Set-Sequence model with Gated Selection
@@
where $\rho(\cdot)$ is an output projection. The gating, with $G_{i,j}\in[0,1]$ allows the summary statistic of unit i to mainly depend on the other units with ’similar’ features to unit i. We recover the standard Set-Sequence model when we set $G_{i,j}=\frac{1}{M}$, for all $i,j$ where $M$ is the number of units. We provide experiments using the Gated Selection mechanism in Section 14.
@@
  • Most importantly, we only compute one G for all times for the sample, this has a large impact on the memory usage, removing the linear scaling with sequence length needed for gradient computations in the selection mechanism.
Appendix D Experiment Details
@@
Table 7: Hyper‑parameter settings for each task.
Hyper‑parameter Synthetic Mortgage Equity
Set‑Seq layers 5 6 6
@@
All the experiments in the paper were conducted on a Linux cluster with 5 NVIDIA RTX A6000 GPUs, each with 49140 MB memory, running on CUDA Version 12.5. The cluster has 256 AMD EPYC 7763 64-Core Processor CPUs. For the synthetic task, each model was trained for less than 1 hour on 1 GPU. For the equities task, each train run (one run per validation year and random seed) took around 30 minutes on a single GPU. The experiments for the mortgage risk prediction task ran for around 1 hour per model training on one GPU.
@@
Appendix E Synthetic Task: Additional Details and Results
E.1 Ablations on the training input units sampling
@@ E.2 Interpretability
E.3 Kalman Filter Baseline
The synthetic task is described in Section 5. We assume that the transition matrix is known, and we also assume that the dynamics of $\lambda_{t}$, $\lambda_{t+1}=\beta\lambda_{t}+\alpha N_{t}$ are known, with the only unknown factor being not observing the true fraction of rare events $N_{t}$, but we only observe $\bar{N}_{t}$, where $\bar{N}_{t}$ is the observed fraction of rare events, only having access to n out of M units. The problem is now to infer the best estimate of $\lambda_{t}$ based on a partial observation of the fraction of rare events. We utilize a Kalman Filter Kalman (1960) to estimate $\lambda_{t}$ with $\hat{\lambda}_{t}$. The resulting filter estimates
@@ E.3 Kalman Filter Baseline
$\hat{\lambda}_{t+1}$ $=\max(\hat{\lambda}_{t+1},0),\quad\text{(to ensure positivity)}$
@@
Appendix F Equity Portfolio Prediction: Additional Details and Results
Figure 8 shows the cumulative returns for the Set-Sequence model compared with the baselines.
Table 9: Out of sample annualized performance metrics net transaction costs 2002 - 2016, trained with a net sharpe ratio objective. Details of the models other than the Set-Sequence model are found in Guijarro-Ordonez et al. (2022). The training setting for the sequence models and Set-Sequence model is in Appendix D.
Model Net Sharpe $\mu$ % $\sigma$ %
@@ Appendix F Equity Portfolio Prediction: Additional Details and Results
We test the Set-Sequence model and the other baseline models under realistic transaction costs. In particular, for each day we subtract the transaction costs from the daily portfolio returns, which are used to compute the Sharpe ratio. Following Guijarro-Ordonez et al. (2022), we use the following model for the transaction cost for portfolio weights $w_{t}$ at time t, and $w_{t-1}$ at t-1:
[…]
The return at time t, can then be expressed as $r_{net,t}=r_{t}-\text{cost}(w_{t},w_{t-1})$. We now train the Set-Sequence model and the Sequence model baselines with a modified objective using the net return Sharpe as the loss function. In Table 9 we show the net Sharpe ratio after costs in Jan. 2002 - Dec. 2016. We see the gain from modeling the cross-section as the Set-Sequence model outperforms all the sequence models in terms of net Sharpe Ratio, as well as the domain specific CNN-Transformer.
@@
Table 10 shows the rank normalized features. In addition to a subset of the features in Chen et al. (2024) we include features for daily return, weekly return, weekly volatility, daily volume. In order to keep the level information, we also include cross-sectional median features in Table 11, following Chen et al. (2024).
Table 10: Firm‑specific characteristics (six categories) used as features in the equity‑portfolio‑optimization task. Construction details are in the Internet Appendix of Chen et al. (2024).
@@
r2_1 Short‑term momentum A2ME Assets to market cap
r12_2 Momentum BEME Book‑to‑market ratio
r12_7 Intermediate momentum C Cash + short‑term inv. / assets
r36_13 Long‑term momentum CF Free cash‑flow / book value
ST_Rev Short‑term reversal CF2P Cash‑flow / price
@@
Rel2High Closeness to 52‑week high
@@
PM Profit margin RF Risk‑free rate
@@
PCM Price‑to‑cost margin
Table 11: Cross‑sectional median firm‑characteristic variables and their stationary transformations ($t\text{Code}$).  Each feature is the cross‑sectional median of the underlying firm characteristic. Note: The transformations ($t\text{Code}$) are (1) no transformation; (2) $\Delta x_{t}$; (3) $\Delta\log(x_{t})$; (4) $\Delta^{2}\log(x_{t})$.
Variable ttCode Variable ttCode Variable ttCode Variable ttCode
@@
Appendix G Mortgage Risk Prediction: Additional Details and Results
G.1 Additional Data Descriptions
We provide additional details about the CoreLogic Loan-Level Market Analytics Dataset. The dataset is over 870 GB and contains month-by-month transition data for mortgages in the US. The dataset starts in 1988, and has data to 2024, with data from 30000 ZIP codes in the US. We restrict ourselves to the top 4 ZIP codes, in terms of number of active loans, in the greater Los Angeles area and leave to follow up work to train on the full dataset. The ZIP Codes are: 92677, with 109642 loans in Orange County; 93065, with 98673 loans in Simi Valley; 91709, with 95497 loans in Chino Hills; 92336 with 94794 loans in Fontana. We follow the same data filtering procedure as in Sadhwani et al. (2020), and use a subset with 52 of the features included there, we filter out loans if any of the following features are not present: FICO score, Original balance, Initial interest rate, and Current State. Some other features we use include Original LTV, Unemployment Rate, Current Interest Rate, and National Mortgage Rate. We deal with missing data in other features with the missing indicator method, see, for example Little & Rubin (2014). After filtering $117,523$ loans remain, each active for on average 45 months, for a total of around 5 million loan month transitions. Figure 10 shows the empirical transition probabilities across the dataset, and Figure 10 shows the transition counts on the sampled test set. Figure 11 shows the number of active loans, the prepayment rate, and the foreclosure rate over time on the dataset. For example, note the elevated foreclosure rates during and after the 2008 financial crisis. Table 12 shows the full set of features for the Set-Sequence model.
@@ G.1 Additional Data Descriptions
Table 12: Feature definitions and possible values. We require that current state, FICO score, Original balance, and Initial Interest Rate are available, and use a missing indicator when other features are unavailable. The feature vector including the missing indicators has dimension 52.
@@
Paid Off, Foreclosure, Real‑Estate Owned
@@
Time since origination <1<1 Year,  <5<5 Years,  <10<10 Years
Original term <17<17 Years
Number of 30‑day delinquencies (past 12 months) 0–12
Number of 60‑day delinquencies (past 12 months) 0–12
Number of 90+‑day delinquencies (past 12 months) 0–12
Number of 30‑day foreclosures (past 12 months) 0–12
@@ G.2 Model Fitting
G.3 Additional Experiments Mortgage Risk Prediction
G.3.1 Interpretability
In Figure 12 in the Appendix, the foreclosure rate, a known source of cross unit dependency, is shown to have a Pearson correlation of 0.67 with the set summary in the first Set-Sequence layer, indicating the interpretability of the learned set summaries.
G.3.2 Transition Analysis
Figure 13 shows the one vs rest AUC conditioned on the initial state for the Set-Sequence model, averaged over 10 seeds of samples on the test set, where for each seed we sample 25 sequences of 2500 loans from the test set.
@@ G.3.2 Transition Analysis
G.3.3 Year-by-year results
@@ Results
We compare the Set-Sequence model, the Gated Selection model, and the models from Sadhwani et al. (2020) in the setting where we do a base training for 15 epochs up until year 2002, and then for each epoch the train-set gets extended one year further. We save a checkpoint each year the model is extended.
In addition we up-weight the more recent train data, this is described in Section G.3.3. Figure 15 and Figure 16 shows the results, comparing the Set-Sequence, Gated Selection (see Appendix C), NN, and Logistic model, all using the same retraining method, and where we evaluate on one year at a time.
Figure 15: Average AUC, Cross-Entropy, and 30dd-60dd AUC for the Set-Sequence, NN, and Logistic model, all using the same retraining method, and where we evaluate on one year at a time.
@@
We see that the challenging years around the financial crisis affects the logistic model the most. However, the relative performance of the Set-Sequence model and the NN model remain broadly the same, although it widens somewhat during 2009.
G.3.4 Interpretability
@@
Appendix H Additional Metric Details
@@
$\operatorname{AUC}(k)$ $=P\!\bigl(y_{1,k}>y_{2,k}\;\big|\,y_{1}=k,\;y_{2}\neq k\bigr)$
$\quad+\frac{1}{2}\,P\!\bigl(y_{1,k}=y_{2,k}\;\big|\,y_{1}=k,\;y_{2}\neq k\bigr).$